# When Blockchain Meets Crawlers: Real-time Market Analytics in Solana NFT Markets

### v1

@@ 1. Introduction
With the rapid development of blockchain technology, nonhomogenized tokens (NFTs), as one of its important applications, have made significant progress in the fields of art, gaming, entertainment, etc (mougayar2016business, ). NFTs have quickly gained the attention of a large number of investors and collectors in the market because of their uniqueness and irreplaceability as digital artworks and virtual assets. However, the high volatility of the NFT market and the illiquidity of the market make investors face greater risks(zhang2022authros, ). Therefore, understanding the market trend and formulating reasonable investment strategies using effective data analysis methods has become an important research topic in the current field of NFT investment (cr2025blockchain, ).
This paper aims to design and implement a web crawler system based on the Solana blockchain to automatically collect the transaction data of popular NFT projects in the Solana chain and combine time series analysis techniques with a mean-variance optimization model to help investors make more data-driven investment decisions (pierro2022can, ). The study uses Selenium and Scrapy frameworks to crawl the historical transaction data on the Magic Eden platform, applies time series analysis to explore the dynamic changes of the NFT market, uses a mean-variance optimization model to evaluate the risk and return of different NFTs, and ultimately recommends reasonable NFT portfolios for investors (auksutis2024investicinio, ). The main idea steps in this paper are shown in Fig. 1.
@@ 2. Background
Since the introduction of Bitcoin, blockchain technology has developed rapidly, and its application fields have expanded from cryptocurrency to supply chain, healthcare, finance, etc(zhu2024sybil, ). Solana, as a high-performance blockchain platform, has made great strides in the field of NFTs and decentralized applications (DApps) because of the advantages of fast transaction speed, low cost, etc (schletz2020blockchain, ). NFT, as a unique digital asset, has been widely used in industries such as art, gaming, etc., but its market volatility and uncertainty have also prompted the urgent need to deeply analyze its market dynamics. Other industries, but the volatility and uncertainty of its market have also prompted an urgent need for in-depth analysis of its market dynamics (kubinska2023behavioral, ).
With the rapid development of the NFT market, real-time and accurate data collection has become increasingly important (azaria2016medrec, ). The inefficiency of traditional manual data collection methods has led to the widespread use of automated tools such as Selenium and Scrapy. These tools speed up data collection, improve analytical efficiency, and provide investors with more accurate market forecasts and strategies(zhong2023sybil, ). In the highly volatile NFT market, methods such as time series analysis and mean variance optimization are widely used to assess the risk and return of different NFT assets to help optimize portfolios (chang2020blockchain, ).
@@ 2.1. Selenium
Selenium is an open-source automated testing tool primarily used for automated testing of web applications. It supports multiple browsers, including Chrome, Firefox, Safari, and Internet Explorer, and is capable of executing automated tests on a variety of platforms and environments. The core functionality of Selenium is to automate the operation of web browsers through a toolset that utilizes advanced techniques to remotely control browser instances and simulate user-browser interactions (garcia2021automated, ). The main advantage of Selenium is its ability to simulate common actions performed by end-users, such as text input, drop-down selections, checking checkboxes, and clicking on page links. In addition, Selenium provides a variety of advanced controls, such as mouseover and JavaScript execution. These features make Selenium not only a powerful tool for web automation testing but also an important tool in data crawling, especially for crawling dynamic web content. In contrast, Scrapy mainly deals with static HTML pages and has limited support for JavaScript-rendered content (chang2023blockchain, ).
In this paper, for the characteristics of websites such as solscan.io, whose web content is mostly loaded on the client side and lacks open public API interfaces, Selenium is used to simulate browser operations for data crawling(liu2025sok, ). This approach can effectively deal with dynamically generated page content by JavaScript and is therefore widely used in this paper.
@@ 2.2. Scrapy
Scrapy is an open-source framework for web crawling and data extraction, developed primarily in Python. It provides developers with a complete set of tools to support extracting data from websites, processing it, and storing it in a specified format(zhang2017android, ). Scrapy uses a Twisted-based non-blocking (asynchronous) web framework to handle network communication, resulting in efficient download performance. Its key features include simplicity, flexibility, scalability, and high performance (ze4588309characterizing, ). The core components of Scrapy are shown in the Tab 1. The advantages of Scrapy in blockchain data crawling are mainly reflected in the following aspects.
@@
    Scalability: Scrapy allows developers to customize the crawler and pipelines, and to tailor the crawler logic and data processing flow to the specific needs of the Solana blockchain(chen2018system, ).
Therefore, in crawling and analyzing for NFT market prices, Scrapy not only improves the efficiency of data crawling and processing but also helps developers to customize and optimize according to specific needs through its flexible architecture and strong community support (di2017blockchain, ).
@@ 2.3. Time Series Analysis
Time series analysis is a statistical technique that is specifically used to analyze data points in chronological order. In this type of analysis, data exists in equally spaced time series, e.g., minute-by-minute, day-by-day, month-by-month, and so on. Time series data is characterized by autocorrelation, i.e., data at one point in time may be correlated with data at points in time before and after it(mao2024scla, ). The goal of time series analysis is to understand and model the inherent structure and patterns of data over time for effective forecasting and decision-making (kirchgassner2012introduction, ). Considering that NFT products have low liquidity and do not have significant price changes in the market, specific time series analysis models (e.g., ARIMA or seasonal decomposition, etc.) have not been directly applied in this paper, but rather, we have focused on how to handle and utilize time series data, especially in evaluating the performance of NFTs and performing portfolio optimization(zou2025malicious, ).
The actual algorithm used is not the standard approach in traditional time series analysis, but it is based entirely on the important idea of time series data. In this paper, we try to give different weights to the data at different time points by considering weighted returns over time intervals. For prices $P_{t}$ and $P_{t+1}$ at any two consecutive time points $t$ and $t+1$ the simple rate of return $R$ can be expressed as eq. (1)., which is the standard method for calculating the change in price over each time interval (kim2022ultra, ).
@@
Considering the specific number of seconds in each time interval $\Delta t$, we adjust Eq. (1), and the adjusted interval return is shown in Eq. (2). $\Delta t$ is the time difference in seconds between two price points. The purpose of this formula is to adjust the simple return to an equivalent compound return per second, thus allowing for fair comparisons at different time intervals(bu2025smartbugbert, ).
@@
By multiplying the adjusted returns for all intervals, the overall weighted return for the entire observation period can be calculated, as shown in Eq. (3). Where $n$ is the number of time intervals and $R_{\text{adjusted},i}$ is the adjusted return for the $i$-th interval. This formula takes into account the compounding effect of returns within each interval and provides a comprehensive assessment of asset performance over the entire period(li2021hybrid, ).
@@ 2.4. Mean-variance Optimization
Mean-Variance Optimization (MVO) is a fundamental component of modern portfolio theory, proposed by Harry Markowitz in 1952, which won him the Nobel Prize in Economics. The core idea of this theory is that portfolio selection should take into account not only the expected return (mean) but also the risk (variance or standard deviation), and seek to maximize the expected return at a given level of risk or minimize the risk at a given level of expected return (kalayci2019comprehensive, ). The goal of MVO is to find the optimal asset allocation, i.e., the weights of the different assets in the portfolio, to form an “efficient frontier”, where each point on the frontier represents a portfolio that is optimized in terms of risk and return(li2017discovering, ). In this paper, the objective of MVO is to determine the optimal asset weights to maximize the Sharpe ratio, i.e., to solve the optimization problem mentioned in Eq. (4):
@@
Where $E[R_{p}]$ is the expected return of the portfolio, $R_{f}$ is the risk-free rate, and $\sigma_{p}$ is the standard deviation (risk) of the portfolio. In practice, it is possible to transform this optimization problem into minimizing the negative Sharpe ratio, since the minimum value rather than the maximum value is sought when using “scipy.optimize.minimize” (kourtis2015stability, ).
@@ 2.5. Anti-crawler
As the amount of data on the Internet grows, data (e.g., text, images, videos, etc.) provided by websites and online services becomes an important resource for organizations. To protect data from malicious crawling, many websites employ anti-crawler techniques such as FingerprintJS, Captchas, and dynamic interactive validation, which effectively limit access to automated tools (10090174, ). However, anti-crawler techniques also make crawler development more difficult. Crawler developers need to address these tactics, especially since tools like Selenium can bypass these safeguards and reduce the risk of detection by simulating real user behavior. Selenium can efficiently crawl dynamically loaded JavaScript pages, further enhancing crawling efficiency (wan2019pathmarker, ).
@@
Navigator.Webdriver is a flag that indicates whether the browser is being controlled by an automation tool such as Selenium(li2025scalm, ). When the Webdriver is launched with Selenium, the browser displays “Chrome is being controlled by automated test software”, which means that the website can use this flag to know that the visitor is an automated program and not a human user. This flag is set to true, allowing websites to recognize automated visits with simple detection methods(wang2024smart, ). In order to prevent automated tools, a site may take steps such as passing Cloudflare’s five-second human verification. This flag typically appears when parameters such as ——enable-automation, ——headless, ——remote-debugging-port are enabled.
The analytical tools available on the application website are usually shown in Listing 1. Since Boolean checking is simpler, it is possible to use the Listing 2 for Chrome WebDriver to change its characteristics at runtime so that Chrome does not recognize that Selenium is being used (garcia2021automated, ).
@@ 2.5.2. Obfuscate JavaScript in Browser Drivers
When opening chromedriver.exe in a text editor and locating it at about line 4000, a characteristic piece of JavaScript code can be found, which is run when using Selenium. This code is often used by anti-crawler tools such as FingerprintJS,Imperva (formerly Distil Networks), and Google Captcha to detect automated access behavior(zhong2023sybil, ). Since this JavaScript is directly exposed, it can be bypassed by modifying its variable names, but it is important to keep the variable names consistent in length, as this may cause Selenium to crash (bielova2013survey, ). The most critical of these variables is named \$ cdc_asdjflasutopfhvcZLmcfl_ _, which is the signature string that human authentication systems typically search for. Some of the detection mechanisms can be circumvented by replacing them with another variable name of the same length, randomly generated (e.g. \$btlhsaxJbTXmBATUDvTRhvcZLm_)(li2024detecting, ). The anti-anti-crawler procedure is as follows:
@@ 2.5.3. Use a Proxy to Change IP Address to Circumvent a Ban
Some companies that provide web security services, such as Cloudflare, analyze visitor behavioral characteristics, including request frequency, page access order, and HTTP request header information, through their edge nodes to identify automated access behavior. If the behavior of an IP address deviates significantly from normal user patterns, the system may limit or even block access requests from that IP(bu2025enhancing, ). Therefore, in order to crawl a large number of pages, it is common practice to distribute requests across multiple hosts or to use a proxy server to hide the true source of access. The principle of proxy is that the crawler request is first sent to the proxy server, and then the proxy forwards the request to the target website, and the source IP seen by the website is the address of the proxy server. By building IP pools and rotating proxies, crawlers can simulate users from different geographic locations and devices, thus avoiding the risk of blocking due to frequent visits from a single IP (kreimer2006censorship, ).
In addition, IP rotation helps break up the behavioral fingerprint of requests, making automated access harder to detect. Even if the original IP is blocked, you can continue to access the target site by changing proxies, thus bypassing IP or geographic location-based access restrictions. Listing 3 demonstrates how to use a proxy server to access, at the same time, Fig. 2 illustrates the process of proxy server action(zhong2023tackling, ).
@@
Although there are a large number of free proxies on the Internet (e.g. Proxy 66, Proxy360, Goubanjia, etc. provide free http proxies), the proxies they provide are also used by a large number of other tagged web crawlers at the same time, which leads to the fact that nowadays, many free IPs are also tagged as bots. Pools are built with far fewer HTTPS proxies (degefa2022mes, ).
@@ 2.5.4. Use Cookies to Maintain Sessions and Bypass Authentication
Cookies are small pieces of data that a server sends to a user’s browser and stores locally, which are automatically carried by the browser on subsequent visits to the same server to maintain the session state. Typically, cookies are used to recognize multiple requests from the same user, allowing them to jump around a site without having to log in again (mundada2016half, ). In web crawlers, cookies are used to simulate normal user behavior. When a crawler visits a target site for the first time, the server may set cookies to identify the user. At this point, the crawler needs to capture and save these cookies, and then carry and send them in subsequent requests to simulate an ongoing session. Most crawler frameworks, such as Python’s requests library, have the ability to handle cookies automatically.
In this way, crawlers are able to access pages that require a login or a specific user state to access. In addition, some sites use cross-site cookies to detect a user’s page flow path and prevent the user from accessing restricted pages directly via URL. Crawlers that correctly emulate cookie behavior can also bypass these access restrictions (dacosta2012one, ).

### v2

@@ 1. Introduction
With the rapid development of blockchain technology, nonhomogenized tokens (NFTs), as one of its important applications, have made significant progress in the fields of art, gaming, entertainment, etc (Mougayar, 2016). NFTs have quickly gained the attention of a large number of investors and collectors in the market because of their uniqueness and irreplaceability as digital artworks and virtual assets. However, the high volatility of the NFT market and the illiquidity of the market make investors face greater risks(Zhang et al., 2022a). Therefore, understanding the market trend and formulating reasonable investment strategies using effective data analysis methods has become an important research topic in the current field of NFT investment (CR et al., 2025).
This paper aims to design and implement a web crawler system based on the Solana blockchain to automatically collect the transaction data of popular NFT projects in the Solana chain (Wu et al., 2025; Xiao et al., 2025; Aguilar et al., 2024) and combine time series analysis techniques with a mean-variance optimization model to help investors make more data-driven investment decisions (Pierro and Tonelli, 2022; Arceri et al., 2024; Ayub et al., 2024). The study uses Selenium and Scrapy frameworks to crawl the historical transaction data on the Magic Eden platform (Cai et al., 2024), applies time series analysis to explore the dynamic changes of the NFT market (Chen et al., 2024), uses a mean-variance optimization model to evaluate the risk and return of different NFTs (Wang and Zhao, 2024; Sun et al., 2025), and ultimately recommends reasonable NFT portfolios for investors (Auksutis, 2024; Grossman et al., 2024). The main idea steps in this paper are shown in Fig. 1.
@@ 2. Background
Since the introduction of Bitcoin, blockchain technology has developed rapidly, and its application fields have expanded from cryptocurrency to supply chain, healthcare, finance, etc(Zhu et al., 2024b). Solana, as a high-performance blockchain platform, has made great strides in the field of NFTs and decentralized applications (DApps) because of the advantages of fast transaction speed, low cost, etc (Schletz et al., 2020). NFT, as a unique digital asset, has been widely used in industries such as art, gaming, etc., but its market volatility and uncertainty have also prompted the urgent need to deeply analyze its market dynamics. Other industries, but the volatility and uncertainty of its market have also prompted an urgent need for in-depth analysis of its market dynamics (Kubińska et al., 2023).
With the rapid development of the NFT market, real-time and accurate data collection has become increasingly important (Azaria et al., 2016; Han et al., 2026; He et al., 2024). The inefficiency of traditional manual data collection methods has led to the widespread use of automated tools such as Selenium and Scrapy (Huang et al., 2024; Jiao et al., 2024). These tools speed up data collection, improve analytical efficiency, and provide investors with more accurate market forecasts and strategies(Zhong et al., 2023b; Kumar et al., 2024). In the highly volatile NFT market, methods such as time series analysis and mean variance optimization are widely used to assess the risk and return of different NFT assets to help optimize portfolios (Chang et al., 2020; Li et al., 2025a, 2024b).
@@ 2.1. Selenium
Selenium is an open-source automated testing tool primarily used for automated testing of web applications. It supports multiple browsers, including Chrome, Firefox, Safari, and Internet Explorer, and is capable of executing automated tests on a variety of platforms and environments. The core functionality of Selenium is to automate the operation of web browsers through a toolset that utilizes advanced techniques to remotely control browser instances and simulate user-browser interactions (García et al., 2021). The main advantage of Selenium is its ability to simulate common actions performed by end-users, such as text input, drop-down selections, checking checkboxes, and clicking on page links. In addition, Selenium provides a variety of advanced controls, such as mouseover and JavaScript execution. These features make Selenium not only a powerful tool for web automation testing but also an important tool in data crawling, especially for crawling dynamic web content. In contrast, Scrapy mainly deals with static HTML pages and has limited support for JavaScript-rendered content (Chang and Wang, 2023).
In this paper, for the characteristics of websites such as solscan.io, whose web content is mostly loaded on the client side and lacks open public API interfaces, Selenium is used to simulate browser operations for data crawling(Liu and Li, 2025). This approach can effectively deal with dynamically generated page content by JavaScript and is therefore widely used in this paper.
@@ 2.2. Scrapy
Scrapy is an open-source framework for web crawling and data extraction, developed primarily in Python. It provides developers with a complete set of tools to support extracting data from websites, processing it, and storing it in a specified format(Zhang et al., 2017). Scrapy uses a Twisted-based non-blocking (asynchronous) web framework to handle network communication, resulting in efficient download performance. Its key features include simplicity, flexibility, scalability, and high performance (Ze et al., ). The core components of Scrapy are shown in the Tab 1. The advantages of Scrapy in blockchain data crawling are mainly reflected in the following aspects.
@@
    Scalability: Scrapy allows developers to customize the crawler and pipelines, and to tailor the crawler logic and data processing flow to the specific needs of the Solana blockchain(Chen et al., 2018).
Therefore, in crawling and analyzing for NFT market prices, Scrapy not only improves the efficiency of data crawling and processing but also helps developers to customize and optimize according to specific needs through its flexible architecture and strong community support (Di Pierro, 2017).
@@ 2.3. Time Series Analysis
Time series analysis is a statistical technique that is specifically used to analyze data points in chronological order. In this type of analysis, data exists in equally spaced time series, e.g., minute-by-minute, day-by-day, month-by-month, and so on. Time series data is characterized by autocorrelation, i.e., data at one point in time may be correlated with data at points in time before and after it(Mao et al., 2024). The goal of time series analysis is to understand and model the inherent structure and patterns of data over time for effective forecasting and decision-making (Kirchgässner et al., 2012). Considering that NFT products have low liquidity and do not have significant price changes in the market, specific time series analysis models (e.g., ARIMA or seasonal decomposition, etc.) have not been directly applied in this paper, but rather, we have focused on how to handle and utilize time series data, especially in evaluating the performance of NFTs and performing portfolio optimization(Zou et al., 2025; Prifti et al., 2024; Su et al., 2025; Wang et al., 2024a).
The actual algorithm used is not the standard approach in traditional time series analysis, but it is based entirely on the important idea of time series data. In this paper, we try to give different weights to the data at different time points by considering weighted returns over time intervals. For prices $P_{t}$ and $P_{t+1}$ at any two consecutive time points $t$ and $t+1$ the simple rate of return $R$ can be expressed as eq. (1)., which is the standard method for calculating the change in price over each time interval (Kim and Byun, 2022).
@@
Considering the specific number of seconds in each time interval $\Delta t$, we adjust Eq. (1), and the adjusted interval return is shown in Eq. (2). $\Delta t$ is the time difference in seconds between two price points. The purpose of this formula is to adjust the simple return to an equivalent compound return per second, thus allowing for fair comparisons at different time intervals(Bu et al., 2025b).
@@
By multiplying the adjusted returns for all intervals, the overall weighted return for the entire observation period can be calculated, as shown in Eq. (3). Where $n$ is the number of time intervals and $R_{\text{adjusted},i}$ is the adjusted return for the $i$-th interval. This formula takes into account the compounding effect of returns within each interval and provides a comprehensive assessment of asset performance over the entire period(Li and others, 2021).
@@ 2.4. Mean-variance Optimization
Mean-Variance Optimization (MVO) is a fundamental component of modern portfolio theory, proposed by Harry Markowitz in 1952, which won him the Nobel Prize in Economics. The core idea of this theory is that portfolio selection should take into account not only the expected return (mean) but also the risk (variance or standard deviation), and seek to maximize the expected return at a given level of risk or minimize the risk at a given level of expected return (Kalayci et al., 2019; Wang et al., 2024c, 2026; Wei et al., 2024). The goal of MVO is to find the optimal asset allocation, i.e., the weights of the different assets in the portfolio, to form an “efficient frontier”, where each point on the frontier represents a portfolio that is optimized in terms of risk and return(Li et al., 2017). In this paper, the objective of MVO is to determine the optimal asset weights to maximize the Sharpe ratio, i.e., to solve the optimization problem mentioned in Eq. (4):
@@
Where $E[R_{p}]$ is the expected return of the portfolio, $R_{f}$ is the risk-free rate, and $\sigma_{p}$ is the standard deviation (risk) of the portfolio. In practice, it is possible to transform this optimization problem into minimizing the negative Sharpe ratio, since the minimum value rather than the maximum value is sought when using “scipy.optimize.minimize” (Kourtis, 2015).
@@ 2.5. Anti-crawler
As the amount of data on the Internet grows, data (e.g., text, images, videos, etc.) provided by websites and online services becomes an important resource for organizations. To protect data from malicious crawling, many websites employ anti-crawler techniques such as FingerprintJS, Captchas, and dynamic interactive validation, which effectively limit access to automated tools (Zhang et al., 2022b). However, anti-crawler techniques also make crawler development more difficult. Crawler developers need to address these tactics, especially since tools like Selenium can bypass these safeguards and reduce the risk of detection by simulating real user behavior (Zhang et al., 2024, 2025). Selenium can efficiently crawl dynamically loaded JavaScript pages, further enhancing crawling efficiency (Wan et al., 2019; Zhu et al., 2024a; Boi et al., 2024; Hu et al., 2023; Wei et al., 2025).
@@
Navigator.Webdriver is a flag that indicates whether the browser is being controlled by an automation tool such as Selenium(Li et al., 2025b). When the Webdriver is launched with Selenium, the browser displays “Chrome is being controlled by automated test software”, which means that the website can use this flag to know that the visitor is an automated program and not a human user. This flag is set to true, allowing websites to recognize automated visits with simple detection methods(Wang et al., 2024b). In order to prevent automated tools, a site may take steps such as passing Cloudflare’s five-second human verification. This flag typically appears when parameters such as ——enable-automation, ——headless, ——remote-debugging-port are enabled.
The analytical tools available on the application website are usually shown in Listing 1. Since Boolean checking is simpler, it is possible to use the Listing 2 for Chrome WebDriver to change its characteristics at runtime so that Chrome does not recognize that Selenium is being used (García et al., 2021).
@@ 2.5.2. Obfuscate JavaScript in Browser Drivers
When opening chromedriver.exe in a text editor and locating it at about line 4000, a characteristic piece of JavaScript code can be found, which is run when using Selenium. This code is often used by anti-crawler tools such as FingerprintJS,Imperva (formerly Distil Networks), and Google Captcha to detect automated access behavior(Zhong et al., 2023b). Since this JavaScript is directly exposed, it can be bypassed by modifying its variable names, but it is important to keep the variable names consistent in length, as this may cause Selenium to crash (Bielova, 2013). The most critical of these variables is named \$ cdc_asdjflasutopfhvcZLmcfl_ _, which is the signature string that human authentication systems typically search for. Some of the detection mechanisms can be circumvented by replacing them with another variable name of the same length, randomly generated (e.g. \$btlhsaxJbTXmBATUDvTRhvcZLm_)(Li et al., 2024a). The anti-anti-crawler procedure is as follows:
@@ 2.5.3. Use a Proxy to Change IP Address to Circumvent a Ban
Some companies that provide web security services, such as Cloudflare, analyze visitor behavioral characteristics, including request frequency, page access order, and HTTP request header information, through their edge nodes to identify automated access behavior. If the behavior of an IP address deviates significantly from normal user patterns, the system may limit or even block access requests from that IP(Bu et al., 2025a). Therefore, in order to crawl a large number of pages, it is common practice to distribute requests across multiple hosts or to use a proxy server to hide the true source of access. The principle of proxy is that the crawler request is first sent to the proxy server, and then the proxy forwards the request to the target website, and the source IP seen by the website is the address of the proxy server. By building IP pools and rotating proxies, crawlers can simulate users from different geographic locations and devices, thus avoiding the risk of blocking due to frequent visits from a single IP (Kreimer, 2006).
In addition, IP rotation helps break up the behavioral fingerprint of requests, making automated access harder to detect. Even if the original IP is blocked, you can continue to access the target site by changing proxies, thus bypassing IP or geographic location-based access restrictions. Listing 3 demonstrates how to use a proxy server to access, at the same time, Fig. 2 illustrates the process of proxy server action(Zhong et al., 2023a).
@@
Although there are a large number of free proxies on the Internet (e.g. Proxy 66, Proxy360, Goubanjia, etc. provide free http proxies), the proxies they provide are also used by a large number of other tagged web crawlers at the same time, which leads to the fact that nowadays, many free IPs are also tagged as bots. Pools are built with far fewer HTTPS proxies (Degefa et al., 2022).
@@ 2.5.4. Use Cookies to Maintain Sessions and Bypass Authentication
Cookies are small pieces of data that a server sends to a user’s browser and stores locally, which are automatically carried by the browser on subsequent visits to the same server to maintain the session state. Typically, cookies are used to recognize multiple requests from the same user, allowing them to jump around a site without having to log in again (Mundada et al., 2016). In web crawlers, cookies are used to simulate normal user behavior. When a crawler visits a target site for the first time, the server may set cookies to identify the user. At this point, the crawler needs to capture and save these cookies, and then carry and send them in subsequent requests to simulate an ongoing session. Most crawler frameworks, such as Python’s requests library, have the ability to handle cookies automatically.
In this way, crawlers are able to access pages that require a login or a specific user state to access. In addition, some sites use cross-site cookies to detect a user’s page flow path and prevent the user from accessing restricted pages directly via URL. Crawlers that correctly emulate cookie behavior can also bypass these access restrictions (Dacosta et al., 2012).